the length of a rectangular playing field is is 5ft less than twice its width.If the perimeter of the playing field is 230ft, find the length and width of the field.
Answer by tutorcecilia(2152)

Use the formula for the perimeter of a rectangle
P=2(Length+width)
P=230
Length=2w-5
Width=w
.
P=2(l+w) [plug-in the values]
230=2[(2w-5)+w] [solve for w]
230=2(3w-5)
230/2=2(3w-5)/2
115=3w-5
115+5=3w-5+5
120=3w
120/3=3w/3
40=w
.
So, if width =40, than the length is:
Length=2w-5=2(40)-5=80-5=75
.
Check by plugging all values back into the original formula and solve:
P=2(length + width)
230=2(75+40)
230=2(115)
230=230 [checks out]